OpenStudy (anonymous):

can somebody please explain how i should solve this? i'd really appreciate it :) Find the indicated arc length. The arc corresponding to a central angle of 35° in a circle of radius 10 feet

hero (hero):

Hint: x = arc length \[\frac{x}{2\pi r} = \frac{35}{360}\]

OpenStudy (anonymous):

r is the radius correct?

hero (hero):

Yes

hero (hero):

In general \[\frac{\text{arc length}}{\text{circumference}} = \frac{\text{arc measure}}{\text{circle measure}}\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so from x/2πr=35/360 how do i continue to solve it

hero (hero):

Isolate x

hero (hero):

of course r = 10 so input that then isolate x

hero (hero):

It's not really difficult. If you know how to solve proportions, you should be able to solve this.

OpenStudy (anonymous):

ok, so what do i use for pi

OpenStudy (anonymous):

3.14?

hero (hero):

3.14 or you can just leave it as \(\pi\)

hero (hero):

It's best to just leave it as \(\pi\)

hero (hero):

Isolate x then reduce the fraction.

OpenStudy (anonymous):

so x/62.83=.097?

OpenStudy (anonymous):

does that work?

hero (hero):

Yes, but it is an approximation, so it's not exact

hero (hero):

First solve the exact, then find the approximate

OpenStudy (anonymous):

how do i solve the exact

hero (hero):

any time you solve a problem, you should post the exact answer, and then after that, post the approximate answer

OpenStudy (anonymous):

i'm not exactly sure how to find the exact, just the approximate :(

hero (hero):

\(\large\frac{x}{2 \pi (10)} = \frac{35}{360}\) \(\large\frac{x}{20\pi} = \frac{35}{360}\) \(\large x = \frac{35}{360} \dot\ 20\pi\) \(\large x = \frac{35 \dot\ 20\pi}{360}\) \(\large x = \frac{35 \pi}{18}\) \(x \approx 6.11\)

hero (hero):

That's how you should solve all math problems that have approximate answers. Post the exact answer first in reduced form. Then post the approximate result.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

it makes much more sense now

OpenStudy (anonymous):

so the arc would be approximately 6.11 degrees?

hero (hero):

ya

OpenStudy (anonymous):

thanks so much you're so great at this stuff

hero (hero):

It's not difficult stuff

OpenStudy (anonymous):

haha well for me it is, but i dont understand it when i read it from a book, this kind of explanation helps me a lot more

OpenStudy (anonymous):

can you please help explain the one i just posted? its the same type of problem, i'd like to get practice with more than one just to be sure that i understand the topic

hero (hero):

Actually, there's a three fraction ratio for this that shows the relationship between length, measure, and area of a circle \[\frac{\text{arc length}}{\text{circumference}} = \frac{\text{arc measure}}{\text{circle measure}} = \frac{\text{area of sector}}{\text{area of circle}}\]

hero (hero):

In the simplest manner: \[\frac{x}{2 \pi r} = \frac{y}{360} = \frac{z}{\pi r^2}\] x = arc length y = arc measure z = area of sector

hero (hero):

the numerator represents a portion of the circle. the demoninator represents the whole circle.

OpenStudy (anonymous):

ohh ok i see

OpenStudy (anonymous):

thanks again :)